Tbone:

Damn quick math question what do y'all get from this equation; 3-3*6+2=

Tbone:

I solved it -3*6=-18 Then I use the rest of the equation(following pemdas) 3-18+2= -16 Drop the three 3-16= -13

Tbone:

But is I solve it straight across the board no pemdas noting I'll get 0*6+2= 0 Drop the two 0+2= 2

TheluvMe:

3-3*6+2 3x6=18 3-18+2=17

TheluvMe:

@tbone wrote:
I solved it -3*6=-18 Then I use the rest of the equation(following pemdas) 3-18+2= -16 Drop the three 3-16= -13
wrong

TheluvMe:

@tbone wrote:
But is I solve it straight across the board no pemdas noting I'll get 0*6+2= 0 Drop the two 0+2= 2
we never do it like this

Tbone:

@theluvme wrote:
@tbone wrote:
I solved it -3*6=-18 Then I use the rest of the equation(following pemdas) 3-18+2= -16 Drop the three 3-16= -13
wrong
This isn't wrong tho according to pemdas

TheluvMe:

@tbone wrote:
@theluvme wrote:
your pemdas is wrong 3-3*6+2 3x6=18 3-18+2 3-18=15 15+2=17 17 is the correct answer
@tbone wrote:
I solved it -3*6=-18 Then I use the rest of the equation(following pemdas) 3-18+2= -16 Drop the three 3-16= -13
wrong
This isn't wrong tho according to pemdas

Tbone:

Pemdas parenthesis expontes multiplication division addition subtraction

TheluvMe:

I know what pemdas is

Tbone:

Then you would know you start with -3*6

TheluvMe:

i did got 18

TheluvMe:

brung it down

TheluvMe:

3-18+2

TheluvMe:

3-18=15

TheluvMe:

15+2=17

Tbone:

When you multiply a negative and a positive the answer is always negative

TheluvMe:

whats the three infront of it

Tbone:

The three Infront of it is a positive

TheluvMe:

okay i apologize

TheluvMe:

the answer is -13

Tbone:

It's cool

Tbone:

Oh okay so the first was right, thanks

TheluvMe:

yep